The convex lens of a copy machine has a focal length of 22.5 cm. A letter to be copied is placed 45.0 cm from the lens.

(a) How far from the lens is the copy paper? cm

(b) How much larger will the copy be? ho
Answer #1

(a) Magnification = image distance / object distance
M = v / u

1 / f = 1 / v + 1 / u

1/v = 1/ f - 1 / u

1 / v = 1/ 22.5 - 1 /45.0 = 0.0444-0.0222 = 0.0222
v=45.04 cm



(b)

M = v /u = 45.04/45 = 1.00089 times larger